# Probing quantum complexity via universal saturation of stabilizer entropies

### v2

@@ 7 Discussion
We find that the SRE converges to the maximum at a critical T-gate density $q_{\text{c},\alpha}$ and time $t_{\text{c},\alpha}$ in the thermodynamic limit. We determine the transition exactly for $\alpha=2$, while for general $\alpha$ we determine the convergence using heuristic models of the Pauli spectrum. For $\alpha=2$, we observe universal behavior around the critical point where the derivative of the SRE can be rescaled onto a single curve for all $N$. This hints that the saturation transition is connected to phase transitions, where universal behavior is commonly found, for example for the transition between different phases of quantum many-body system [Osterloh et al.(2002)Osterloh, Amico, Falci, and Fazio] or at complexity transitions in classical and quantum algorithms [Zhang et al.(2022)Zhang, Sone, and Zhuang].
@@
We find that random Hamiltonian evolution has a fundamental separation in nonstabilizerness complexity: Simulating random evolution using Clifford-based algorithms to arbitrary precision is already hard for very small times $t\gtrsim 1/\text{poly}(N)$. In contrast, up to $t=\text{const}$ one can efficiently certify the fidelity of the evolved state within some fixed error $\Delta F\approx 4t$. We show the algorithm in Appendix D by certifying in respect to a stabilizer approximation of the state. In contrast, for Cliffords doped with T-gates, simulation and certification complexity correlate, and become intractable for the same T-gate density $q\gtrsim\log(N)$.
@@ Acknowledgements
Appendix A GUE evolution and random basis evolution
@@
Appendix B Min-relative entropy of magic scaling for random evolution
@@
Appendix C SRE, min-relative entropy and robustness
@@
Refer to captionaRefer to caption Refer to captionaRefer to caption
Figure S3: Relationship of $\alpha$-SRE with min-relative entropy $D_{\text{min}}$ and log-free robustness LR. Here, we rescaled $D_{\text{min}}$ and LR to pose lower bounds on SRE. a) Clifford+T circuit for $N=4$ qubits against T-gate density $q$. b) Evolution with random Hamiltonian sampled from GUE against time $t$.
Appendix D State certification via Pauli measurements and SREs
@@
For (approximate) Pauli fidelity estimation, where one checks the fidelity against the closest stabilizer state, the sampling complexity is related to the closest stabilizer state. As such, one can approximate the fidelity efficiently as long as the $\alpha=2$ SRE $M_{2}$ is sufficiently small. This holds true even when $\alpha<1$ SREs has become extensive.

### v3

@@ 7 Discussion
We find that the SRE converges to the maximum at a critical T-gate density $q_{\text{c},\alpha}$ and time $t_{\text{c},\alpha}$ in the thermodynamic limit. We determine the transition exactly for $\alpha=2$, while for general $\alpha$ we determine the convergence using heuristic models of the Pauli spectrum. For $\alpha=2$, we observe universal behavior around the critical point where the derivative of the SRE can be rescaled onto a single curve for all $N$. This hints that the saturation transition is connected to phase transitions, where universal behavior is commonly found, for example for the transition between different phases of quantum many-body system [Osterloh et al.(2002)Osterloh, Amico, Falci, and Fazio] or at complexity transitions in classical and quantum algorithms [Zhang et al.(2022)Zhang, Sone, and Zhuang]. While we can show the universality analytically for $\alpha=2$, a follow-up work observes strong indications of universality numerically for $\alpha=1$ by studying system sizes up to $N=24$ [Tarabunga and Haug(2025)]. We believe universality holds for all $\alpha$, which would be interesting to study in future works. For example, using recent methods for Clifford circuits doped with magic, it may be possible to find analytic expressions beyond $\alpha=2$ [Leone et al.(2025)Leone, Oliviero, Hamma, Eisert, and Bittel].
@@
We find that random Hamiltonian evolution has a fundamental separation in nonstabilizerness complexity: Simulating random evolution using Clifford-based algorithms to arbitrary precision is already hard for very small times $t\gtrsim 1/\text{poly}(N)$. In contrast, up to $t=\text{const}$ one can efficiently certify the fidelity of the evolved state within some fixed error $\Delta F\approx 4t$. We show the algorithm in Appendix F by certifying in respect to a stabilizer approximation of the state. In contrast, for Cliffords doped with T-gates, simulation and certification complexity correlate, and become intractable for the same T-gate density $q\gtrsim\log(N)$.
@@ Acknowledgements
Appendix A Comparison maximal magic and Haar random states
In the main text, we derive the maximal SRE $M_{\alpha}^{\text{max}}$ as function of $\alpha$. To derive this result, we approximate it with the SRE of a state that has effectively a completely uniform Pauli spectrum with $\bra{\psi}P\ket{\psi}=2^{-N}$ for $P\neq I$.
We now provide an alternative way to compute the maximal magic. It is known that Haar random states have in the limit of large $N$ the maximal possible magic [Liu and Winter(2022)]. The average SRE of Haar random states has been derived analytically in Ref. [Turkeshi et al.(2025)Turkeshi, Dymarsky, and Sierant] in the limit of large $N$ as
$M_{\alpha}^{\text{Haar}}=\frac{1}{1-\alpha}\ln\left[\frac{(4^{N}-1)(2(2^{N-1}+1)^{-1})^{\alpha}\Gamma(\alpha+\frac{1}{2})}{\sqrt{\pi}2^{N}}+2^{-N}\right]\,,$ (S1)
where $\Gamma(x)$ is the Gamma function. Now, we look at the asymptotics for large $N$, where we find
$M_{\alpha}^{\text{Haar}}\approx\frac{1}{1-\alpha}\ln\left[\frac{2^{N}2^{-N\alpha}\Gamma(\alpha+\frac{1}{2})}{\sqrt{\pi}}+2^{-N}\right]\,.$ (S2)
Now, let us regard $\alpha>2$. Here, the second term dominates and we have
$M_{\alpha>2}^{\text{Haar}}\approx\frac{1}{1-\alpha}\ln\left[2^{-N}\right]=-\frac{1}{1-\alpha}N\ln 2\equiv M_{\alpha>2}^{\text{uniform}}\,,$ (S3)
matching the maximal magic we derived via the flat Pauli spectrum. Next, for $\alpha<2$, the first term dominates with
$M_{\alpha<2}^{\text{Haar}}\approx\frac{1}{1-\alpha}\ln\left[\frac{2^{N(1-\alpha)}\Gamma(\alpha+\frac{1}{2})}{\sqrt{\pi}}\right]=N\ln 2+\frac{1}{1-\alpha}\ln\left[\frac{\Gamma(\alpha+\frac{1}{2})}{\sqrt{\pi}}\right]\equiv M_{\alpha<2}^{\text{uniform}}+\text{const}\,,$ (S4)
matching the flat Pauli spectrum result up to constant corrections.
Appendix B GUE evolution and random basis evolution
@@
Appendix C Min-relative entropy of magic scaling for random evolution
@@
Appendix D Further numerical results
Here, we show additional numerical results on the time evolution with random Hamiltonians.
First, in Fig. S3 we show SRE $M_{\alpha}(t)$ against time $t$ for different $\alpha$, where the dashed lines are the approximations $M_{\alpha}^{\text{approx}}$ derived in the main text. Here, we simulate the dynamics using statevector simulation. In Fig. S3b, we show the error $|M_{\alpha}-M_{\alpha}^{\text{approx}}|/M_{\alpha}$ relative to the approximations, where we find a close match for $\alpha>1$ for all $t$, while $\alpha>1$ matches only for $t<1$. Note that the approximations were derived for small $t$, yet we find that they work quite well even for larger $t$.
Refer to captionaRefer to caption Refer to captionbRefer to caption
Figure S3: $\alpha$-SREs of random Hamiltonian evolution. a) SRE $M_{\alpha}(t)$ against time $t$ for $N=12$ and different $\alpha$, where dashed lines are approximations $m_{\alpha}^{\text{approx}}$ of Eq. (32), Eq. (29), Eq. (30) and Eq. (31). The sudden change of the dashed line for $\alpha<1$ is due to the change from Eq. (29) to Eq. (30) at $12^{N(1-\alpha)}(1-F^{2})^{\alpha}=$ 1. b) Relative error $|M_{\alpha}-M_{\alpha}^{\text{approx}}|/M_{\alpha}$ against $t$.
In Fig. S4, we show the SRE $M_{\alpha}(t)$ against qubit number $N$ at different times $t$ under random Hamiltonian evolution. We find that the error depends only weakly on $N$, which gives evidence that our approximations are accurate even for larger $N$.
Refer to captionaRefer to caption Refer to captionbRefer to caption
Refer to captioncRefer to caption Refer to captiondRefer to caption
Refer to captioneRefer to caption Refer to captionfRefer to caption
Figure S4: $\alpha$-SREs of random Hamiltonian evolution against $N$ for different times a,b) $t=10^{-2}$, c,d) $t=10^{-1}$, and e,f) $t=1$. a,c,e) SRE $M_{\alpha}(t)$ where dashed line is fit with approximations Eq. (32), Eq. (29), Eq. (30) and Eq. (31). b,d,f) Relative error $|M_{\alpha}-M_{\alpha}^{\text{approx}}|/M_{\alpha}$ to approximations $M_{\alpha}^{\text{approx}}(t)$.
Appendix E SRE, min-relative entropy and robustness
@@
Refer to captionaRefer to caption Refer to captionbRefer to caption
Figure S5: Relationship of $\alpha$-SRE with min-relative entropy $D_{\text{min}}$ and log-free robustness LR. Here, we rescaled $D_{\text{min}}$ such that it poses an upper bound on SRE for $\alpha=2$, and LR as upper bound for $\alpha\geq 1/2$. a) Clifford+T circuit for $N=4$ qubits against T-gate density $q$. b) Evolution with random Hamiltonian sampled from GUE against time $t$.
Appendix F State certification via Pauli measurements and SREs
@@
For (approximate) Pauli fidelity estimation, where one checks the fidelity against the closest stabilizer state, the sampling complexity is related to the closest stabilizer state. As such, one can approximate the fidelity efficiently as long as $\alpha=2$ SRE $M_{2}$ is sufficiently small. This holds true even when the corresponding SRE with $\alpha<1$ SREs have already become extensive.